# Formalization of Auslander–Buchsbaum–Serre criterion in Lean4

### v2

@@ 5.4.1. Auslander–Buchsbaum–Serre Criterion
With the results in Subsections  5.2 and  5.3, we can finally state and proof the Auslander–Buchsbaum–Serre criterion, following the original version in [1], as follows.W
@@ 5.4.2. Regular Ring
A ring $R$ is a regular ring if $R_{\mathfrak{p}}$ is regular local ring for every prime ideal $\mathfrak{p}$.
@@
localization_isRegular : p : Ideal R, (_ : p.IsPrime),
IsRegularLocalRing (Localization.AtPrime p)
@@
A ring $R$ is regular if and only if $R_{\mathfrak{m}}$ is regular local ring for every maximal ideal $\mathfrak{m}$.
@@ 5.4.3. Hilbert’s Syzygy Theorem
theorem Polynomial.isRegularRing_of_isRegularRing [IsRegularRing R] :
IsRegularRing R[X]
@@ 6.1. Remarks About Implementation
In the formalization of category theory in Lean4, one often needs to carefully consider universe levels. For example, in this project we must specify the universe in which the $\operatorname{Ext}$ functor lives. In our application, we mainly use the instance for having $\operatorname{Ext}$ for categories with enough projectives W. For the category of $R$-modules in universe $v$ with $R$ in universe $u$, if we assume $R$ is $v$-small (denoted [Small.{v} R]), then we can obtain $\operatorname{Ext}$ in any universe greater than or equal to $v$. In our implementation, we only explicitly specify the universe level of $\operatorname{Ext}$ when it actually appears in the statement of a theorem; otherwise we hide it or fix a universe level. For example, for projectiveDimension and globalDimension, we use the standard universe following previous definitionsW. For the definition of depth, with $R$ in universe u and assuming Small.{v} R, in ModuleCat.{v} R we consider $\operatorname{Ext}$ in universe v, which is the smallest universe in which we obtain $\operatorname{Ext}$ using CategoryTheory.hasExt_of_enoughProjectives W. None of this matters much in practice, since there are already definitions for changing universes W, which simply pass through shifted hom; moreover, if the category is $R$-linear, then this is an $R$-linear equivalence and the additional structure is preserved.
The remaining nuisance is the universe of the modules on which we take $\operatorname{Ext}$. Since the commutation of $\operatorname{Ext}$ and Ulift is not yet fully formalized yet, some definitions still partially depend on a choice of universe, such as moduleDepth, projectiveDimension and globalDimension. Currently we only now moduleDepth is invariant of universe in some good conditions by using the regular sequence characterization. However, as we know that $\operatorname{Ext}$ and Ulift do commute, it would be easy to prove that moduleDepth and projectiveDimension is invariant under arbitrary isomorphisms once that fact is formalized. As for globalDimension, from Lemma  2.13 we know that it suffices to consider the supremum over finitely generated modules. Notice that when assuming Small.{v} R and Small.{v’} R, there is a correspondence between finitely generated modules in ModuleCat.{v} R and ModuleCat.{v’} R. The globalDimension would be invariant when assuming the ring is small.
@@ 6.3. Future Works
In Subsection  5.4.3, we proved that the global dimension of $k[x_{1},\dots,x_{n}]$ is $n$, which is only slightly stronger than the statement that every finitely generated module over $k[x_{1},\dots,x_{n}]$ has a projective resolution of length nn—the weakest version of Hilbert’s Syzygy Theorem. These projective resolutions can indeed be upgraded to free resolutions using the Quillen–Suslin theorem (see [18][23]) or other constructions specific to polynomial rings over fields (see [8, Corollary 19.8]). Additionally, there is a graded version of this theorem (see [8, Corollary 19.7]); however, with the current setup, it may not be feasible to formalize that version.

### v3

@@ 5.4.1. Auslander–Buchsbaum–Serre Criterion
With the results in Subsections  5.2 and  5.3, we can finally state and prove the Auslander–Buchsbaum–Serre criterion, following the original version in [1], as follows.W
@@ 5.4.2. Regular Ring
A ring $R$ is a regular ring if $R_{\mathfrak{p}}$ is a regular local ring for every prime ideal $\mathfrak{p}$.
@@
localization_isRegular : p : Ideal R, (_ : p.IsPrime), IsRegularLocalRing (Localization.AtPrime p)
@@
A ring $R$ is regular if and only if $R_{\mathfrak{m}}$ is a regular local ring for every maximal ideal $\mathfrak{m}$.
@@ 5.4.3. Hilbert’s Syzygy Theorem
theorem Polynomial.isRegularRing_of_isRegularRing [IsRegularRing R] : IsRegularRing R[X]
@@ 6.1. Remarks About Implementation
In the formalization of category theory in Lean4, one often needs to carefully consider universe levels, for example the universe level of $\operatorname{Ext}$ groups. In our implementation, for constructions on general abelian categories such as projectiveDimension and globalDimension, we use the standard universe level or leave the HasExt instance as parameters following previous definitionsW. When working on the category of $R$-modules in universe $v$, if we assume $R$ is $v$-small (denoted [Small.{v} R]), the category has enough projectives, which allows us to obtain $\operatorname{Ext}$ in universe level vvW. This is also the only universe level of $\operatorname{Ext}$ we use throughout the development. The reason is that users should not have to specify the universe parameter when considering $\operatorname{Ext}$ types, so a given category should have at most one HasExt instance. In this setup, the universe level v is therefore the only reasonable choice.
The remaining nuisance is the universe of the modules on which we take $\operatorname{Ext}$. Since the commutation of $\operatorname{Ext}$ and Ulift is not yet fully formalized yet, some definitions still partially depend on a choice of universe, such as moduleDepth, projectiveDimension and globalDimension. Currently we only know moduleDepth is invariant of universe in some good conditions by using the regular sequence characterization. However, as we know that $\operatorname{Ext}$ and Ulift do commute (up to isomorphism), it would be easy to prove that moduleDepth and projectiveDimension is invariant under arbitrary linear isomorphisms once that fact is formalized. As for globalDimension, from Lemma  2.13 we know that it suffices to consider the supremum over finitely generated modules. Notice that when assuming Small.{v} R and Small.{v’} R, there is a correspondence between finitely generated modules in ModuleCat.{v} R and ModuleCat.{v’} R. The globalDimension would be invariant when assuming the ring is small.
@@ 6.3. Future Works
In Subsection  5.4.3, we proved that the global dimension of $k[x_{1},\dots,x_{n}]$ is $n$, which is only slightly stronger than the statement that every finitely generated module over $k[x_{1},\dots,x_{n}]$ has a projective resolution of length nn—the weakest version of Hilbert’s Syzygy Theorem. These projective resolutions can indeed be upgraded to free resolutions using the Quillen–Suslin theorem (see [18][23]) or other constructions specific to polynomial rings over fields (see [8, Corollary 19.8]). Additionally, there is a graded version of this theorem (see [8, Corollary 19.7]); however, with the current setup, it may not be feasible to formalize the graded version.